Today’s shift goal was to check the alignment of the HWS DET probe beam, in relation to the previously observed issues (59347, 59385, 59387).
Once the morning activity was completed, around 15.50 UTC Piernicola went to the DET lab to first of all change the HWS with the CCD. See Fig. 1 for references of EDB HWS optics.
With the ITF locked in NI Single Bounce we firstly see a beam near to the edge of the camera (Fig. 2 - left) and, once we move the coming back beam acting on EDB_Har_M5 by ~3 round horizontally, clipped (Fig. 2 right)
At this point we realized that we were looking at the probe beam coming from INJ. By switching OFF INJ sled and ON DET sled, we saw a good beam: Piernicola reversed the action on EDB_Har_M5 (but not completely, he did ~1.5 round back) to center the HWS DET probe beam coming back from the NI TM (see Fig. 3 for the comparison wrt the last picture taken on Nov. 28th, 2022 (57924)):
The beam shape is good, no sign of clipping is visible, and seemed analogous to the last one.
At this point we made some tests:
-
When CITF was locked, the beam seemed to move slightly, but remained on the camera sensor [not sure if the illumination level changed because of alignment or due to a slightly different sled constant value; from now on it is 1.05 and will not be changed until the end of the test] → Fig. 4.
-
Trying to reach CARM NULL, we managed to see the beam at CARM NULL 3F, before an unlock: the beam remained inside the sensor even if it moved toward the bottom-center of the camera, with still good shape, illumination counting passed from ~1750 to ~1600.
-
Setting the NI single bounce the beam came back to condition in Fig. 3; if in NI single bounce the NE mirror is aligned, the illumination change significantly (~1200 → ~1900) and the beam seemed to move toward the bottom-right; indeed, probably the back reflection of the DET probe beam coming from the NE is strong enough to dominate the camera illumination (see Fig. 5)
-
We try to check also the effect of tilting the NI MIR (the optic that seemed to "move more" during the Gentle Unlock); the movements (reported in Fig. 6) results in a slight change of the illumination (~1900 → ~1800)
At the end, we sat back the NI single bounce + SR aligned, checked the last time the CCD image and replaced the CCD with the HWS (illumination visible in Fig. 7).
The HWS DET probe beam seemed to properly propagate toward the NI TM and to get back from it.
Some issues (59430) prevented the CARM NULL. If the CARM NULL is recovered, a NI absorption automatic measurement will be launched to see if the slight adjustment performed on the coming back beam on top of the CCD/HWS sensor improves the measurement results.